OpenStudy (anonymous):

a right triangle has a short leg of 67 cm. and the longer leg has a measurement of 122 cm. how long will be the hypotenuse of the triangle?

OpenStudy (anonymous):

use a^2+b^2=c^2

OpenStudy (anonymous):

139,18692467326088601081441877375

OpenStudy (anonymous):

no 67^2+122^2=19373, c=sqrt19373